OpenStudy (anonymous):

Will FAN The ages of students in a school are normally distributed with a mean of 15 years and a standard deviation of 2 years. Approximately what percent of the students are between 14 and 18 years old?

OpenStudy (anonymous):

A. 24.17% B. 62.47% C. 30.85% D. 93.32%

OpenStudy (anonymous):

Well, we have : mean = 15 years SD = 2 years let be F the distribution of the N(0,1) is the z-table, then P( 14 < Age < 18) = P( - 0.5 < (Age - mean)/SD < 1.5) = F(1.5) - F( -0.5) = .93319 - .30854 -----------------------> P( 14 < Age < 18) = .62465 second answer 2. classical data to know by heart : m = 3 , σ = .25 P( 2.5 < Time < 3.5) = P( Normal distribution € [ m - 2σ , m + 2σ] ) # 0.95

OpenStudy (ajanijones):

You will find its, "0.4332" and for the left side, its "0.1915" and their sum is "0.6247". Just multiply with "100" to turn the probability in to percentage. The answer is "62.47%"

OpenStudy (ajanijones):

Your answer is B. Hope I helped out.

OpenStudy (anonymous):

i was getting there

OpenStudy (anonymous):

Thanks a lot!!! I really appreciate it! both of you!

OpenStudy (ajanijones):

@your_mom017 Here's Medal to you do deserve it. :)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

It was correct for future users!